undeadknight26 (undeadknight26):

Simplify: 7 square root of 3 end root minus 4 square root of 6 end root plus square root of 48 end root minus square root of 54

jigglypuff314 (jigglypuff314):

if you mean \[7\sqrt{3}- 4\sqrt{6} + \sqrt{48} - \sqrt{54}\] then think of it this way \[7\sqrt{3} - 4\sqrt{6} + \sqrt{16* 3} - \sqrt{9*6}\rightarrow 7\sqrt{3} - 4\sqrt{6} + \sqrt{16}*\sqrt{3} - \sqrt{9}*\sqrt{6}\] so you get 7√3 - 4√6 + 4√3 - 3√6 and just like how you can only add x variable with x variables √3 can only be added to values with √3 and so on... so 7√3 - 4√6 + 4√3 - 3√6 = (7√3 + 4√3) + (- 4√6 - 3√6) 11√3 - 7√6

undeadknight26 (undeadknight26):

is this right? Is the simplified form of 2square root of 3 • square root of 12 rational? Yes No****

jigglypuff314 (jigglypuff314):

the answer is yes because 2√3 * √12= 2√3 * √4*3 = 2√3 * √4*√3= 2√3 * 2√3 and know that when numbers in ratical are same and are multiplied gets you just number like √3 * √3 = √(3*3) = √9 = 3 so 2√3 * 2√3 = 4 * 9 = 36

jigglypuff314 (jigglypuff314):

sorry I meant 2√3 * 2√3 = 4 * 3 = 12